OpenStudy (blehh10):

(x+1/x^2-25) * (x+5/x^2+8x+7) Please explain... x+1/(x+5)(x+7) 1/(x+5)(x+7) 1/(x-5)(x+7) 1/(x-5)(x-7)

jimthompson5910 (jim_thompson5910):

What would x^2-25 factor to? hint: difference of squares

OpenStudy (blehh10):

(x+5)(x-5)

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what would x^2+8x+7 factor to? hint: list out all the ways to multiply to 7. Then find out which pair of factors add to 8

OpenStudy (blehh10):

I already did this I just don't understand how to put it together.

jimthompson5910 (jim_thompson5910):

tell me what x^2+8x+7 factors to

OpenStudy (blehh10):

(x+1)(x+7)

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{x+1}{x^2-25}*\frac{x+5}{x^2+8x+7}\] turns into \[\Large \frac{x+1}{(x+5)(x-5)}*\frac{x+5}{(x+1)(x+7)}\]

OpenStudy (blehh10):

So... what does that mean?

jimthompson5910 (jim_thompson5910):

Multiply the terms straight across to combine the fractions making \[\Large \frac{x+1}{(x+5)(x-5)}*\frac{x+5}{(x+1)(x+7)}\] turn into \[\Large \frac{(x+1)(x+5)}{(x+5)(x-5)(x+1)(x+7)}\]

jimthompson5910 (jim_thompson5910):

Notice these (x+1) terms pair up. One in the numerator. One in the denominator \[\Large \frac{{\color{red}{(x+1)}}(x+5)}{(x+5)(x-5){\color{red}{(x+1)}}(x+7)}\]

jimthompson5910 (jim_thompson5910):

The (x+1) terms will cancel since x/x = 1 (where x is nonzero) \[\Large \frac{{\color{red}{\cancel{(x+1)}}}(x+5)}{(x+5)(x-5){\color{red}{\cancel{(x+1)}}}(x+7)}\]

OpenStudy (blehh10):

Ohh. Thanks I get it now. I forgot about the cancelling!

OpenStudy (blehh10):

You also cancel (x+5)

jimthompson5910 (jim_thompson5910):

After that cancellation of the (x+1) terms, we're left with this \[\Large \frac{x+5}{(x+5)(x-5)(x+7)}\] I'm sure you see at this point the (x+5) terms cancel as well

jimthompson5910 (jim_thompson5910):

`You also cancel (x+5)` yes correct

OpenStudy (blehh10):

So since you helped me can I give you a medal or something? I don't know if I can?

jimthompson5910 (jim_thompson5910):

to give out a medal, click on the blue "best answer" button

OpenStudy (blehh10):

Where is that?

OpenStudy (adrimit):

the upper right corner of the comment you want to give the medal to

OpenStudy (blehh10):

So I did it?

OpenStudy (adrimit):

yes

OpenStudy (blehh10):

Okay thanks, I haven't really used openstudy.

jimthompson5910 (jim_thompson5910):

yes @blehh10 you gave out the medal, thank you

OpenStudy (blehh10):

Really, thank you!

OpenStudy (adrimit):

welcome to open study :)

OpenStudy (blehh10):

Thanks!